# Gravitational wave probes of particle dark matter: a review

### v3

@@ VI.3.3 Constraints on luminous dark photons clouds
As alluded to in Section V, if the dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it could produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen et al. (2023a). This additional coupling implies that soon after exponential growth of the boson cloud begins, electromagnetic fields act on charged particles in the vicinity of the BH, producing electron/positron pairs essentially out of vacuum, i.e. through the photon-assisted Schwinger mechanism Sauter (1931); Heisenberg and Euler (1936); Schwinger (1951). At this point, a pair-production cascade ensues Dunne et al. (2009); Monin and Voloshin (2010), sourced by the electromagnetic fields, creating a plasma around the BH in a state of turbulent quasi-equilibrium. This plasma emits copious amounts of electromagnetic radiation, sourced by energy injections through dissipative processes in the cloud (e.g. at magnetic reconnection sites, Landau damping, turbulence), potentially even periodically. The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and it could be periodic Siemonsen et al. (2023a). If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to, and potentially indistinguishable from, known pulsars in the ATNF catalog Manchester et al. (2005).
@@
The lack of observation of CWs from pulsars, and the potential of dark-photon clouds to act as “pulsar mimickers”, motivates a search for such clouds around targets in the ATNF catalog that could be mislabeled as pulsars. In Ref. Mirasola et al. (2025), the authors consider this possibilty by analyzing a subset of, spinning up “pulsars” from the ATNF catalog that might actually be but stellar-mass BHs with a dark-photon luminous cloud around them. Using a variety of methods – one that assumes perfect phase coherence between the observed electromagnetic radiation and GW emission, one that allows small deviations in frequency around a narrow band of the observed frequency of each compact object, and one that is most robust against theoretical uncertainties in the signal but least sensitive to GW emission –, upper limits on the dark photon/ordinary photon kinetic mixing parameter are set as a function of boson mass for the 34 sources considered in this analysis, as shown in Fig. 22. To obtain this plot, particular BH populations and electromagnetic emissions are assumed to determine the number of events on the color bar, so the bounds are not necessarily robust against changes in these choices. For the BH population, the relevant distributions are over position in the galaxy (black holes follow the stellar distribution) and mass (Salpeter), spin (uniform) and age (uniform) of the BH systems.
Though particular population models are assumed, the constraints only require ten “events” to have been observable, which is conservative with respect to the potentially $\mathcal{O}(10^{3})$ that would be expected at such coupling/mass combinations, as shown in the right-hand panel of Fig. 3 of Mirasola et al. (2025). Here, “events” just means the number of rotating BHs that could have given off a detectable signal, that is, with a GW amplitude larger than the upper limit strain amplitude, in an assumed population of $10^{8}$ BHs in the galaxy. In the end, a dark photon mass range between, roughly $[10^{-13},10^{-12}]$ eV could not have a kinetic mixing more than roughly $[10^{-9},10^{-7}]$.
@@
Figure 22: Taken from Mirasola et al. (2025). For each of the targets of the search in Mirasola et al. (2025), the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits. The color corresponds to the number of events that should have been detected, assuming a “pessimistic” population model for BHs. The bounds reproduced from other experiments can be found in O’Hare (2020), specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi et al. (2009); Caputo et al. (2020) (dark gray), and from measurements by Planck and unWISE galaxies of cross correlations between temperature anisotropies in the CMB McCarthy et al. (2024) (light gray).
@@ VI.3.4 Stochastic gravitational-wave background constraints
The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers. In Tsukada et al. (2019), the authors designed a model of the stochastic GW background arising from isolated BHs formed in two ways: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars. The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background Tsukada et al. (2019). No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence. The major reason for the inability to robustly exclude boson masses is that there exist degeneracies between the boson mass and BH astrophysical quantities, such as spins. The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al. (2021), in particular disfavoring vector boson cloud masses between $[0.8,6]\times 10^{-13}$ eV at 95% confidence. Additionally, Ref. Yuan et al. (2022) has also disfavored bosons with masses between $[1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors. Furthermore, the stochastic GW background from tensor bosons around rotating BHs has been computed and searched for in Guo et al. (2024b), and has disfavored boson masses between $[10^{-13.4},10{-11.7}]$ eV depending on the range of uniform spin distributions assumed Guo et al. (2024b).
@@ VI.4 Prospects for GW probes of boson clouds
The chosen targets of GW searches should also be evaluated. While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki et al. (2015a), these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi et al. (2019), though could be seen up to Gpc now if they are composed of vector bosons Jones et al. (2023). On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen and East (2020), and so the robustness of search results by targeting remnants would have to be evaluated.
@@
All-sky searches for such boson cloud systems could also be enhanced. At the moment, especially in light of boson self-interactions that could alter the GW signal significantly. Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al. (2023, 2025), could be useful in a variety of contexts. Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider any grid in spin-up, which limited it to be sensitive to boson clouds whose frequency drift did not exceed one frequency bin during the entire observation time. Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal.
@@
It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions. Ref. Hoof et al. (2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions. The authors encourage that posteriors on mass and spin be made publicly available in order to incorporate into their future analyses. Furthermore, Ref. Guo et al. (2024c) shows that bosons could be transferred from a growing level to a depleting one if both BHs in a binary system have clouds around them, which may affect the GW waveform and hence would need to be considered in future GW searches.
VII GW probes of soliton DM
If DM has a mass of $\sim 10^{-24}-10^{-22}$ eV Hui et al. (2017), its Compton wavelength will be comparable to the size of the galactic halo, which actually limits the minimum mass of DM Sin (1994). In this mass regime DM could form so-called “solitons”, self-gravitating structures prevented from collapsing from degeneracy pressure, at the centers of galaxies Schive et al. (2014). To probe this kind of DM, these masses imply that the oscillation frequencies will be of $\mathcal{O}(\text{nHz})$, making pulsar timing arrays the ideal probe Aghaie et al. (2024). In essence, DM would affect the waveform arising from the superposition of inspiraling supermassive BHs, causing the BHs to experience dynamical friction, thus weaking the signal strain in the nHz regime. But this is not the only way to detect this mass regime of ultralight dark matter: as we will show, LISA and next-generation ground-based GW interferometers could also be used. To probe this kind of DM with conventional GW sources, low-frequency information needs to be contained in relatively higher-frequency carrier GW. The GW frequency becomes modified during its propagation between its source and the earth as it passes through the DM soliton, i.e. it is “heterodyned”, or gravitationally redshifted, analogously to the Sachs-Wolfe effect that induces fluctuations in the cosmic microwave background Sachs and Wolfe (1967).
VII.1 Impact on phase evolution of binary systems
Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo et al. (2016); Babak et al. (2017). The long-durations of such signals permit the possibility of detecting a low-frequency modulation when passing through the ultralight dark matter soliton Wang and Zhong (2023); Brax et al. (2024); Blas et al. (2025). Ref Brax et al. (2024) has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future. Unfortunately, even at these masses, the density of DM would have to be at least $10^{4}$ times higher than in the solar system in order to observe any changes to the phase evolution GW. This is broadly consistent with previous work by Wang and Zhong (2023), which assumed an enhancement of the DM density by eight orders of magnitude relative to that on earth, and showed that DM with a mass of $\sim 10^{-23}$ eV could be detected with LISA. Furthermore, Ref. Blas et al. (2025) assumes a representative population of white dwarf binaries, a fraction of which are detectable in LISA, and concludes that DM masses above $\sim 10^{-23}$ eV will remain unconstrained by LISA. However, the mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV could be constrained if DM has stronger, quadratic couplings to ordinary matter.
VII.2 Imprint of DM on CW signal from neutron star
In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce changes in the CWs arising from deformed neutron stars, thus allowing the possibility of detecting this effect in future ground-based detectors Blas et al. (2025). These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles (2017). In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more neutron stars spinning down due to the emission of CWs Pagliaro et al. (2023), all of which could be modulated by ultralight dark matter solitons.
Using CWs from deformed neutron stars, the soliton mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV becomes accessible in ET for both linear and quadratic couplings, and improves with the number of neutron stars that are detected. The constraints will even outperform existing pulsar timing array ones, since the signal-to-noise ratio scales linearly with the carrier GW frequency. This is shown in LABEL:fig:lambda12res for LISA, ET and Cosmic Explorer, and demonstrates the powerful probe that CWs from neutron stars will be of this particular DM model.
VII.3 Pulsar-timing constraints on soliton ultralight dark matter
@@
The evidence for a GW background from NANOGrav results can be used to put constraints on soliton DM Aghaie et al. (2024). Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit, which suppresses the GW strain at nanohertz frequencies. The supermassive BHs are too massive to be prevented from merging, as would be the case for lighter BHs Hui et al. (2017); Bar-Or et al. (2019). Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig. 1 of Afzal et al. (2023), Ref. Aghaie et al. (2024) derives constraints on the soltion DM mass-to-BH mass ratio parameter space, as shown in Fig. 24. These bounds are derived by using Eq. 17 of Aghaie et al. (2024), which depends on how much energy is lost via GW emission normally and in the presence of the ultralight dark matter soliton, and on the differential density populations of supermassive BHs. The latter can be written analytically in terms of the the galaxy-stellar mass functions, galaxy merger rates and galaxy pair fractions Chen et al. (2019). It is then a matter of integrating Eq. 17 in Aghaie et al. (2024) over these different distributions (which may depend on decently uncertain astrophysics) for different choices of ultralight dark matter mass to derive the constraint on mass ratio of soliton DM to supermassive BH mass, and determine if the integrated strain exceeds the error bar on the GW strain obtained by NANOGrav.
@@
Figure 24: Taken from Aghaie et al. (2024). Upper limits from NANOGrav data on the DM mass and solition-BH mass ratio parameter space, obtained by noting that inspiraling supermassive BH background for which NANOGrav found evidence would be modulated by the dynamical friction induced by soliton DM. Magenta lines show the scaling relations derived through DM-only numerical simulations (solid), and ones in which supermassive BHs dominate scaling relations suggested by DM-only numerical simulations and SMBH dominance (dashed).

### v4

@@ VI.3.3 Constraints on luminous dark photon clouds
As alluded to in Section V, if dark photon DM kinetically mixes with the ordinary photon and forms clouds around rotating BHs, it would produce both electromagnetic and GW signals, which, if detected, could enable a major multi-messenger discovery Siemonsen et al. (2023a). This additional coupling implies that soon after exponential growth of the boson cloud begins, electromagnetic fields act on charged particles in the vicinity of the BH, producing electron/positron pairs essentially out of vacuum, i.e. through the photon-assisted Schwinger mechanism Sauter (1931); Heisenberg and Euler (1936); Schwinger (1951). At this point, a pair-production cascade ensues Dunne et al. (2009); Monin and Voloshin (2010), sourced by the electromagnetic fields, creating a plasma around the BH in a state of turbulent quasi-equilibrium. This plasma emits copious amounts of electromagnetic radiation, sourced by energy injections through dissipative processes in the cloud (e.g. at magnetic reconnection sites, Landau damping, turbulence). The luminosity of these sources depends on the strength of the kinetic mixing parameter, but could be at the level of supernovae or known pulsars, and and also periodic Siemonsen et al. (2023a). If the dark photon has a mass around $\sim 10^{-12}$ eV, it will induce the formation of clouds around stellar-mass BHs, and give off radio emission that would look remarkably similar to that from known pulsars in the ATNF catalog Manchester et al. (2005).
@@
The absence of observed CWs from pulsars, combined with the possibility that dark-photon clouds could act as “pulsar mimickers,” motivates searches for such clouds around sources in the ATNF catalog that might be misidentified as pulsars. Ref. Mirasola et al. (2025) explores this scenario by analyzing a subset of spinning-up “pulsars” in the catalog that could instead be stellar-mass BHs surrounded by luminous dark-photon clouds. Three analysis strategies are considered: one assuming perfect phase coherence between the electromagnetic emission and any associated GW signal; one allowing for small frequency deviations around the measured pulsar frequency; and one that is most robust to theoretical uncertainties but least sensitive to GW emission. From these methods, the authors derive upper limits on the dark-photon–photon kinetic mixing parameter as a function of boson mass for the 34 sources studied, shown in Fig. 21.
To produce these bounds, assumptions are made about both the BH population and their electromagnetic signatures, so the precise results are not fully robust against variations in these choices. The adopted BH distributions include a stellar-tracing spatial distribution, a Salpeter mass function, uniform spin and age distributions, and a total galactic population of $10^{8}$ BHs. Even under these assumptions, the analysis requires only $\sim 10$ detectable “events” (i.e., BHs with GW amplitudes exceeding the strain upper limit), which is conservative compared to the $\mathcal{O}(10^{3})$ events expected at certain coupling–mass values (see Fig. 3 of Mirasola et al. (2025)). Ultimately, the study excludes dark-photon masses in the range $\sim[10^{-13},10^{-12}]$ eV with kinetic mixing larger than roughly $10^{-9},10^{-7}]$.
@@
Figure 21: Taken from Mirasola et al. (2025). For each of the radio pulsar mimickers targeted in Mirasola et al. (2025), the upper limit at 95% confidence on the kinetic mixing parameter $\epsilon$ is plotted as a function of the boson mass for which at least 10 systems would have been detected with signal amplitudes above the strain upper limits. Each vertical line corresponds to one radio pulsar, whose frequency fixes the mass of the dark photon. The color corresponds to the number of events that should have been detected, assuming a “optimistic” population model for BHs, i.e. using a larger maximum mass and spin for the assumed BH population with respect to the “pessimistic” population also constrained. The bounds reproduced from other experiments can be found in O’Hare (2020), specifically from the conversion of CMB photons into dark photons, which would cause distortions in the black body spectrum measured by COBE/FIRAS Mirizzi et al. (2009); Caputo et al. (2020) (dark gray), and from measurements by Planck and unWISE galaxies of cross correlations between temperature anisotropies in the CMB McCarthy et al. (2024) (light gray). The upper limits arising from this search surpass those from existing experiments, and indicate that between $10$ and $10^{4}$ events could have been detected at couplings of between $\sim[10^{-10},10^{-7}]$.
@@ VI.3.4 Stochastic gravitational-wave background constraints
The superposition of GWs arising from annihilating scalar boson clouds could form a stochastic GW background detectable by current-generation GW interferometers. In Tsukada et al. (2019), the authors designed two models of the source of this stochastic GW background: (1) from remnants of core collapse supernovae, and (2) from remnants of mergers of BHs or neutron stars. The authors also performed search on LIGO O1 data using cross-correlation to look for evidence of a stochastic GW background from scalar boson clouds Tsukada et al. (2019). No signal was found, and no range of boson masses could be robustly excluded unfortunately, except under optimistic assumptions about BH formation rates and spin distributions, which would disfavor the regime $2\times 10^{-13}\leq m_{b}\leq 3.8\times 10^{-13}$ eV at 95% confidence. Here, “optimistic’ means assuming rapidly spinning and frequently forming BHs in the galaxy. This limitation highlighted the degeneracy between boson mass and poorly constrained astrophysical parameters, particularly BH spins.
The search was repeated in the second observing run as well, finding no evidence for any stochastic GW background of vector boson clouds around rotating BHs Tsukada et al. (2021), in particular disfavoring vector boson cloud masses between $0.8,6]\times 10^{-13}$ eV at 95% confidence. Ref. Yuan et al. (2022) has also disfavored bosons with masses between $1.5,17]\times 10^{-13}$ eV depending upon what is assumed for the uniform spin distribution, with wider constraints present when higher-spinning BHs dominate the priors.
Beyond scalar and vector cases, Ref. Guo et al. (2024b) investigated tensor bosons, computing the corresponding stochastic GW background and performing a dedicated search. Their analysis disfavored boson masses between $10^{-13.4}$ and $10^{-11.7}$ eV, again with the precise range depending on the assumed range of uniform spin distribution .
Taken together, these studies demonstrate that the stochastic GW background provides a powerful probe of ultralight scalar, vector, and tensor clouds, but the strength of the resulting constraints depends strongly on assumptions about black-hole formation, mass, and spin distributions. This mirrors the caveats discussed in Section VI.3.1, where similar assumptions underlie the derived limits. Moving forward, it is therefore essential to develop robust and consistent methods for setting constraints that minimize reliance on specific black-hole population parameters while retaining astrophysical interpretability.
VI.3.5 Constraints from the spins of detected binary black holes
We note that constraints on boson clouds can be inferred from the spin measurements of binary black hole systems Arvanitaki et al. (2015a, 2017); Brito et al. (2017); Baryakhtar et al. (2017); Cardoso et al. (2018); Ghosh and Sachdeva (2021). Recently, a Bayesian method has been developed by combining the information from each system Ng et al. (2021b), assuming a distribution for the initial BH spins and requiring that the BHs have had enough time to undergo superradiance before merging. Such a study has been applied to the second GW transient catalog (GWTC-2) Abbott et al. (2021b), and has resulted in essentially a joint posterior on the scalar boson mass and the spin distribution of BHs at their formation Ng et al. (2021a). The latter is required because BHs born with high spins that have been spun down by superradiance is partially degenerate with BHs simply born with low spins in a universe without superradiance. The study presented in Ng et al. (2021a) tentatively rules out scalar bosons in the mass range $1.3,2.7]\times 10^{-13}$ eV.
These constraints from measured BH spins can disfavor particular masses or even confirm the presence of an ultralight boson; however, they rely on assumptions about the spin distribution at formation, on whether the boson clouds have time to form (which could be impacted by eccentricity at formation Peters (1964); Wen and Gair (2005)), and are a function of the (strongest) BH mergers that we see with GWs.
Thus, these kinds of analyses complement those previously that look for CWs around isolated, rotating BHs. Both have their advantages: CW analyses described above allow us to probe unknown sources and known, nearby sources whose signals we expect to be stronger than those arising from remnants of mergers. Likewise, the spins and masses of BHs formed in binary system are well measured, and the constraints obtained in Ng et al. (2021a) will only improve as we detect more and more BHs.
@@ VI.4 Prospects for GW probes of boson clouds
The chosen targets of GW searches should also be evaluated. While having accurate measurements of the mass, spin and age of remnants of binary black hole mergers would allow us to ignore astrophysical population models Arvanitaki et al. (2015b), these systems are currently too far to be detected if the clouds are composed of scalar bosons Isi et al. (2019), though could be seen up to Gpc now if they are composed of vector bosons Jones et al. (2023). On the flipside, GWs from vector boson clouds suffer from even more theoretical uncertainties than scalars Siemonsen and East (2020), and so the robustness of search results by targeting remnants would have to be evaluated.
@@
All-sky searches for such boson cloud systems could also be enhanced, especially in light of boson self-interactions that could alter the GW signal significantly as discussed in Section V.3. Potentially, new methods to track quickly-evolving waveforms, such as those in Jones et al. (2023, 2025), could be useful in a variety of contexts. Furthermore, the all-sky search specifically for scalar boson cloud systems did not consider the possibility that the cloud could spin-up, which, in practice, fixed a maximum spin-up to which the search was sensitive (such that the spin-up did not induce a freqeuncy shift by more than one bin $\delta f$ in $T_{\text{obs}}$, i.e. $\dot{f}T_{\text{obs}}<\delta f\equiv 1/T_{\text{FFT}}$). Such modifications to searches, coupled with additional theoretical understanding of the GW signal itself, could increase chances of detecting a signal by exploring a wider parameter space and allowing the cloud to contract over time.
@@
It is also difficult to say with certainty that a particular boson mass is “ruled out”, since the constraints are either population agonistic or depend on some assumptions about BH formation, spin or mass distributions. Ref. Hoof et al. (2024) attempts to unify the way in which constraints are set in a Bayesian way by handling the subtleties associated with population assumptions. The authors in Ref. Hoof et al. (2024) encourage that posteriors on mass and spin, obtained through analyses of X-ray data from BHs, be made publicly available in order to incorporate into their future work.
VII GW probes of soliton dark matter
Ultralight DM in the mass range $\sim[10^{-24},10^{-22}]$ eV has a Compton wavelength comparable to the size of galactic halos Hui et al. (2017); Sin (1994). In this regime, DM can form self-gravitating structures, or solitons, at the centers of galaxies, which remain stable due to quantum degeneracy pressure Schive et al. (2014).
A natural approach to probing this mass range of DM is through pulsar timing arrays, whose frequency sensitivity aligns with the nanohertz oscillation frequencies expected from DM. In this case, solitons can exert dynamical friction on binary systems, accelerating their inspiral and suppressing the gravitational wave strain amplitude at nanohertz frequencies Aghaie et al. (2024).
Alternatively, soliton DM can be framed as a low-frequency modulation of relatively higher-frequency GW. The GW frequency is modified during its propagation through the DM halo, undergoing a gravitational redshift or ’heterodyning’, analogous to the Sachs-Wolfe effect that imprints low-frequency modulations on high-frequency photons in the cosmic microwave background Sachs and Wolfe (1967). Two CW sources have been considered as high-frequency carriers for this modulation Blas et al. (2025): (1) inspiraling white-dwarf binaries, observable by space-based GW interferometers, and (2) non-axisymmetric rotating neutron stars that spin down over time.
In what follows, we explore how these different probes can be used to constrain the properties of soliton. In Section VII.1, we demonstrate how data from pulsar timing arrays can be used to set constraints on soliton DM. We then explore two other approaches in Section VII.2 and Section VII.3: detecting CWs from white-dwarf binaries and isolated neutron stars, which could reveal nanohertz modulations induced by soliton DM.
VII.1 Constraints from pulsar timing arrays
@@
The evidence for a GW background from NANOGrav results can be used to constrain soliton DM Aghaie et al. (2024). Essentially, ultralight dark matter would induce dynamical friction in the inspiral of supermassive BHs, speeding up the orbit and suppressing the GW strain at nanohertz frequencies. Using the most accurate limit at 3.92 nHz on the GW strain from supermassive BHs in the right-hand panel of Fig. 1 of Afzal et al. (2023), Ref. Aghaie et al. (2024) derives constraints on the soliton DM mass-to-BH mass ratio parameter space, as shown in Fig. 22. These constraints are obtained using Eq. 17 from Aghaie et al. (2024), which models energy loss due to GW emission in the presence of ultralight dark matter. This equation incorporates astrophysical inputs such as galaxy-stellar mass functions, galaxy merger rates, and galaxy pair fractions Chen et al. (2019). The constraints are obtained by integrating Eq. 17 over these distributions, taking into account variations in the assumptions about the ultralight dark-matter mass and the population of supermassive BHs. However, these constraints are sensitive to assumptions regarding astrophysical parameters. For example, the specific values chosen for galaxy merger rates, stellar mass functions, and the fraction of galaxies with supermassive BHs affect the final upper limits.
@@
Figure 22: Taken from Aghaie et al. (2024). Upper limits on the DM mass and soliton-to-supermassive black hole mass ratio parameter space, derived from NANOGrav data. These constraints are based on the notion that the GW background from the inspiral of supermassive BHs, for which NANOGrav has found evidence, would be modulated by the dynamical friction exerted by soliton dark matter. Specifically, the presence of soliton DM alters the supermassive BH inspiral dynamics, resulting in a suppression of the GW strain at nanohertz frequencies. The magenta lines represent two distinct sets of scaling relations derived from numerical simulations: solid lines correspond to DM-only simulations, where the soliton DM is the sole interacting component, while dashed lines incorporate the influence of supermassive BHs, suggesting a modification of the scaling relations due to the presence of these massive objects.
VII.2 Dynamical friction on white dwarf binaries
Space-based GW detectors will observe the quasi-infinite inspiral of galactic white dwarf binaries and extreme-mass ratio inspirals of ordinary compact objects with much heavier BHs Bartolo et al. (2016); Babak et al. (2017). The long-durations of such signals permit the possibility of detecting a low-frequency modulation when passing through the ultralight dark matter soliton Wang and Zhong (2023); Brax et al. (2024); Blas et al. (2025). Ref. Brax et al. (2024) has found that the impact of dynamical friction on the phase evolution of GW signals from white dwarf binaries exceeds that of the low-frequency DM-induced modulation for masses above $10^{-21}$ eV, essentially providing two orders of magnitude of mass that could be probed in the future. Unfortunately, even at these masses, the density of DM would have to be at least $10^{4}$ times higher than in the solar system in order to observe any changes to the phase evolution GW. This is broadly consistent with previous work by Wang and Zhong (2023), which assumed an enhancement of the DM density by eight orders of magnitude relative to that on earth, and showed that DM with a mass of $\sim 10^{-23}$ eV could be detected with LISA. Furthermore, Ref. Blas et al. (2025) assumes a representative population of white dwarf binaries, a fraction of which are detectable in LISA, and concludes that DM masses above $\sim 10^{-23}$ eV will remain unconstrained by LISA. However, the mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV could be constrained if DM has stronger, quadratic couplings to ordinary matter. Again, the possibility that DM couples to the standard model in some way is able to enhance the signal strength, as found in Sections II and IV for different types of DM.
VII.3 Nanohertz modulations of continuous gravitational waves from neutron stars
In addition to modulating the millihertz frequency of GWs from binaries in space-based GW detectors, soliton ultralight dark matter could induce modulations in the rotational frequency of neutron stars in a way distinct from ordinary spinning down, deformed neutron stars, thus affecting CW signals and allowing the possibility of detecting this effect in future ground-based detectors Blas et al. (2025). These GWs are sourced from the rotational power of the neutron stars, thus spinning down the neutron stars as GWs are emitted Riles (2017). In ET and Cosmic Explorer, we expect to detect $\mathcal{O}(100)$ or more canonical neutron stars spinning down due to the emission of CWs Pagliaro et al. (2023), which could be modulated by ultralight dark matter solitons.
Using CWs from deformed neutron stars, the soliton mass regime $\sim[2\times 10^{-22},3\times 10^{-21}]$ eV becomes accessible in ET for both linear and quadratic couplings, and improves with the number of neutron stars that are detected. The constraints will even outperform existing pulsar timing array ones, since the SNR scales linearly with the carrier GW frequency. This is shown in Fig. 23 for LISA, ET and Cosmic Explorer, and demonstrates the powerful probe that CWs from neutron stars will be of this particular DM model.
Figure 23: Taken from Blas et al. (2025). Projected constraints on the linear (left) and quartic coupling parameters (right) for both LISA and ET from low-frequency soliton DM modulating (“heterodyning”) high-frequency carrier GW signals via changes to the moment of inertia of neutron stars, which alters CWs from deformed neutron stars (ET) and the GWs from inspiraling binary black holes (LISA). The curves for ET assume one or 200 detections of CWs from isolated neutron stars. The LISA curve is determined by the impact of soliton DM on GW frequencies from binary systems. The olive regions are excluded by the pulsar timing array analysis in Smarra et al. (2024), while blue and red are constrained away via Cassini tests of general relativity (violet) and on a stochastic GW background (red).
We note that in this case, DM leaves a particular imprint on the GW signal itself, arising from deformed, isolated neutron stars. This differs from the discussion in Section II, where DM interacts directly with the components of the interferometer, causing differential strains that are distinct from the expected signal from GWs. In contrast, here we are looking for a deviation in a canonical GW source due to DM, rather than an interaction with the interferometer itself.
Furthermore, though the effect described in Section II.3.2 of dilatons interacting with neutron stars is similar to that discussed here, it arises from a different physical mechanism. While dilatons alter the fundamental constants via a new coupling that leads to a oscillatory moment of inertia of neutron stars, solitons interact directly with the internal structure of the neutron star. This difference leads to solitons modulating the GW frequency, while dilatons would alter the pulse time of arrival in pulsar timing array experiments.
In summary, solitons could significantly impact the GW signals arising from supermassive black hole binaries, white-dwarf binaries, and asymmetrically rotating neutron stars. These signals, which are observed using pulsar timing arrays, space-based interferometers, and ground-based interferometers, respectively, can be modulated by the presence of soliton dark matter. For supermassive BH binaries, solitons can induce dynamical friction on the orbit, leading to detectable changes in the nanohertz frequency range. White-dwarf binaries, could show low-frequency modulations in their GW emission, potentially revealing the imprint of soliton DM on the millihertz carrier signals. Similarly, asymmetrically rotating neutron stars, whose CWs are monitored by ground-based detectors, could exhibit phase shifts by passing through soliton DM. These potential signatures across multiple observational platforms make soliton DM a promising candidate for future studies in GW astronomy, and provide interesting probes of the DM landscape.